Sweet Harmony Crafts has determined that when x hundred Dobros are built, the average cost per Dobro can be estimated by C(x)=.1x^2-.7x+2.425 where C(x) is in hundreds of dollars. What is the minimum average cost per Dobro and how many Dobros should be built to achieve that minimum?

Answer by stanbon(75887)

Sweet Harmony Crafts has determined that when x hundred Dobros are built, the average cost per Dobro can be estimated by C(x)=.1x^2-.7x+2.425 where C(x) is in hundreds of dollars. What is the minimum average cost per Dobro and how many Dobros should be built to achieve that minimum?
-----------------------------------
C(x)=.1x^2-.7x+2.425 where C(x) is in hundreds of dollars.
---------
Minimum occurs when x = -b/2a = -[-0.7/(2*0.1)] = 3.5 hundred
----------
C(3500) = 1.22 million hundred = 1.22*10^8 = 122,000,000 [# of Dobros]
---------------------
Cheers,
Stan H.